# Invariants of persistence modules defined by order-embeddings

### v2

@@ Abstract.
One of the main objectives of topological data analysis is the study of discrete invariants for persistence modules, in particular when dealing with multiparameter persistence modules. In many cases, the invariants studied for these non-totally ordered posets $\mathcal{P}$ can be obtained from restricting a given module to a subposet $\mathcal{X}$ of $\mathcal{P}$ that is totally ordered (or more generally, of finite representation type), and then computing the barcode (or the general direct sum decomposition) over $\mathcal{X}$.
We consider in this paper general order-preserving embeddings of representation-finite subposets $\mathcal{X}$ into $\mathcal{P}$ and study systematically the invariants obtained by decomposing the restriction of a given $\mathcal{P}$-module $M$ to $\mathcal{X}$ into its indecomposable summands. The restriction functor from $\operatorname{\mathrm{mod}}\mathcal{P}$ to $\operatorname{\mathrm{mod}}\mathcal{X}$ is well-studied, and it is known to be exact and admits both left and right adjoint functors, known as induction and co-induction functors. This allows us to obtain new homological insights, and also to re-interpret previous results. We use this approach also to determine bases of the image of these invariants, thus generalizing the concept of signed barcodes which is considered in the literature in relation to stability results.
It turns out that considering only order-embeddings of one fixed poset $\mathcal{X}$ into the poset $\mathcal{P}$, and studying the set of all indecomposables obtained from $\mathcal{X}$ introduces a lot of redundancy. We therefore also study iterated embeddings of several posets of increasing sizes, while limiting attention to only some indecomposables (that have not been obtained from embedding of smaller posets previously).
@@ 1.1. Setting
It turns out that, in many cases, these invariants studied for non-totally ordered posets $\mathcal{P}$ can be obtained from restricting a given module to a subposet $\mathcal{X}$ of $\mathcal{P}$ that is totally ordered (or more generally, of finite representation type), and then computing the barcode (or the general direct sum decomposition) over $\mathcal{X}$. This approach has been described in e.g. [AENY23b, DKM24, HNOX, La18, LW] introducing certain subposets $\mathcal{X}$ having the form of linear subposets, “zigzags”, or “tours” and “courses”, all referring to a poset whose indecomposable representations can be described via intervals. (The full classification of those finite posets whose indecomposable representations can all be described via intervals can be found in [Cha69] and [AET25, Section 5].)
@@
To fix notation, consider an order-embedding $f:\mathcal{X}\to\mathcal{P}$; that is, a poset morphism that is an isomorphism onto its image. The restriction of a $\mathcal{P}$-module $M$ to $f(\mathcal{X})$ can be described completely by counting the multiplicities of indecomposable direct summands. We assume $\mathcal{X}$ to be representation-finite, and thus obtain a discrete invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{f}$ of rank $|\operatorname{\mathrm{Ind}}\mathcal{X}|$, the number of indecomposable $\mathcal{X}$-modules. Repeating this process for all possible embeddings $f:\mathcal{X}\to\mathcal{P}$ and collecting the result into a vector indexed over all embeddings, we obtain the invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ describing all multiplicities of indecomposable $\mathcal{X}$-modules obtained from all possible ways to restrict the $\mathcal{P}$-module $M$ to a subposet of form $\mathcal{X}$. One also defines an equivalent invariant $\mathbf{dimh}_{\mathcal{X},\mathcal{P}}$ by calculating the Hom-dimensions from the set of indecomposable $\mathcal{X}$-modules to the restrictions of the persistence module along all order embeddings. See Definition 3.3 and Theorem 3.6 for details.
Consider for example the poset $\mathcal{X}_{1}$ with one element. Then $\mathcal{X}_{1}$ has only one indecomposable representation, and its multiplicity in a general module is given by the dimension. Thus listing all possible embeddings of $\mathcal{X}_{1}$ into $\mathcal{P}$ and counting multiplicities amounts to describing the dimension vector, so $\mathbf{mult}_{\mathcal{X}_{1},\mathcal{P}}=\operatorname{\mathbf{dim}}_{\mathcal{P}}$. Consider now the poset $\mathcal{X}_{2}=\{1\to 2\}$. This poset is representation-finite, and the multiplicity of the indecomposable representation given by the interval $[1,2]$ can be computed by calculating the rank of the map $M_{1}\to M_{2}$. This allows one to show that the multiplicity invariant $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}$ is equivalent to the well-studied rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}}$, see Corollary 3.8. More generally, as discussed in Example 3.5, the “generalized rank invariant” of [KM21] and the “compressed multiplicities” of [AENY23b] both record the multiplicity of the sincere interval modules $\mathbb{I}_{\mathcal{X}}$ over certain subsets $\mathcal{X}\subseteq\mathcal{P}$.
@@ 1.2. Results
For several subposets $\mathcal{X}$, we address the question of how the multiplicities in $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ can be computed in terms of linear algebra. See Examples 3.7 and 3.15 and the calculations in Section 3. In each case, the multiplicties of indecomposable summands of $M$ can be obtained computing ranks or intersections given by the structure maps of the module $M$. More generally, we show in Theorem 3.6 that the invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ can be obtained by calculating the Hom-dimensions from the set of indecomposable representations to the given persistence module, thus reducing the problem to solving a system of linear equations.
Besides computability, the second and equally important quality of an invariant for persistence modules is the stability. For the rank invariant, this question has been addressed in [La18], and further explored in [BOO24, BOOS24] where the authors single out a class of special persistence modules, the “hook modules”, which are defined as indicator modules of certain shapes called hooks: For $a<b$ in $\mathcal{P}$, let $[a,b[\ =\{c\in\mathcal{P}\mid a\leq c\not\geq b\}$. The interval module $\mathbb{I}_{[a,b[}$ is called a hook module. Note that $\mathbb{I}_{[a,b[}$ is the cokernel of the natural map $\mathbb{I}_{[b,\infty)}\rightarrow\mathbb{I}_{[a,\infty)}$ between indecomposable projective $\mathcal{P}$-modules. It is shown in [BOO24] that the hook modules play the role of projective objects in a relative homology theory given by some exact structure $\mathcal{E}$, thus every persistence module $M$ admits a resolution by hook modules. One can then show that these hooks form a basis for the (image of the) rank invariant (see Definition 4.11), which in turn allows one to define a “signed barcode”, leading to a stability result for the considered invariant [BOOS24].
The viewpoint of relative homological algebra has further been extended in many recent works (see e.g. [AENY23a, BBH24, BBH, BOO24, BOOS24, CGRST24, OS24]), but we are only now able to conceptually explain the appearance and importance of hook modules and their generalizations: Studying embeddings $f:\mathcal{X}\to\mathcal{P}$, the corresponding restriction functor $\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ admits both a left adjoint $f_{!}:\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$ and a right adjoint $f^{*}:\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$. One can then use the adjunction formula to show that the set
=$\mathcal{H}=\{f_{!}U\mid U\mbox{ indecomposable }\mathcal{X}\mbox{-module and }f:\mathcal{X}\to\mathcal{P}\mbox{ an embedding}\}$
@@
splits for all embeddings $f:\mathcal{X}\to\mathcal{P}$. See Proposition 4.8. Coming back to the rank invariant, we study embeddings of the poset $\mathcal{X}_{2}$, and since every indecomposable $\mathcal{X}_{2}$-module is the cokernel of a map between indecomposable projectives, we see easily that the set $\mathcal{H}$ above is formed exactly by the hook modules in this case.
@@
$f(1)$$f(2)$$f(3)$$k$$\mathcal{X}=\mathcal{X}^{\prime}_{3}$$f(3)$$f(2)$$f(1)$$k$$k$$k$$k^{2}$$\mathcal{X}=\mathcal{X}^{\prime\prime}_{3}$
Already in [BOO24], they specified a second basis for the rank invariant, given by rectangle modules $\mathbb{I}_{[a,b]}$ with support $[a,b]$, which are somewhat easier to deal with than the hook modules (note that one needs to consider right-open rectangles when the poset $\mathcal{P}$ is infinite). We aim to generalise this result here in terms of general embeddings: Given a poset embedding $f:\mathcal{X}\rightarrow\mathcal{P}$ and $U\in\operatorname{\mathrm{mod}}\mathcal{X}$, the adjunction formulas yield a natural map $\theta_{f}U:f_{!}U\rightarrow f_{*}U$. We denote $\Theta_{f}U=\mathrm{Im}(\theta_{f}U)$. (This is the so-called “intermediate extension functor”, see [AET25, Section 4.2] for further details and historical context.) One easily verifies that in case $\mathcal{X}=\mathcal{X}_{2}$, the set
=$\mathcal{R}=\{\Theta_{f}U\mid U\mbox{ indecomposable }\mathcal{X}\mbox{-moduleand }f:\mathcal{X}\to\mathcal{P}\mbox{ an embedding }\}$
@@
=$\mathcal{H}=\{f_{!}U\mid U\mbox{ indecomposable }\mathcal{X}\mbox{-module and }f:\mathcal{X}\to\mathcal{P}\mbox{ an embedding}\}$
produces the same $\mathcal{P}$-modules several times from different embeddings $f$ and different indecomposables $U$, for instance the simple modules are double counted whenever the images of two different embeddings $f_{1},f_{2}:\mathcal{X}\to\mathcal{P}$ intersect. It is therefore more efficient to first collect all simple $\mathcal{P}$-modules (once) via embeddings of the one-point poset $\mathcal{X}_{1}$, then consider all two-dimensional $\mathcal{P}$-modules obtained from the interval $[1,2]$ by embedding the poset $\mathcal{X}_{2}=\{1\to 2\}$ into $\mathcal{P}$ and so on. In other words, instead of studying the set of all indecomposables obtained from embeddings of one poset $\mathcal{X}$, we study now embeddings of several posets of increasing sizes, but limit attention to only some indecomposables (that have not been obtained from embedding of smaller posets previously). The setting we consider here is where we fix a family of posets =$\mathcal{F}=\{\mathcal{X}_{i}\}$ together with certain embeddings into $\mathcal{P}$, and then we define $\mathbf{mult}_{\mathcal{F}}$ to count the multiplicity of the characteristic indecomposable module $\mathbb{I}_{\mathcal{X}_{i}}$ in the restriction of $M$ to $\mathcal{X}_{i}$ along the various embeddings. So only one indecomposable is considered for each poset embedding. As discussed in Example 5.8, both the generalized rank invariant of [KM21] and the various compressed multiplicities of [AENY23b] can be realized in the form $\mathbf{mult}_{\mathcal{F}}$.
@@
With :=$\mathcal{F}:=\{\mathcal{X}_{1},\mathcal{X}_{2}\}$ for $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ the one- and two-point poset as before, we can recover the result that the set $\mathcal{R}$ (the rectangles in this case) forms a basis for $\mathbf{mult}_{\mathcal{F}}$.
Theorem 1.1 (Theorem 5.15).
@@
=$\mathcal{R}=\{\Theta_{f}(U)\mid(f,U)\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}\times\operatorname{\mathrm{Ind}}\mathcal{X}\}$
@@ 2. $\mathbb{Z}$-linear invariants
An object $M\in\operatorname{\mathrm{mod}}\mathcal{P}$ is also called a $\mathcal{P}$-module. We denote by $M_{a}$ the vector space $M(a)$, and by $M_{a\to b}$ the element in $\operatorname{\mathrm{Hom}}_{k}(M_{a},M_{b})$ given by the functor $M$ applied to the pair $a\leq b\in\mathcal{P}$ We further denote $\operatorname{\mathbf{dim}}_{a}M:=\dim_{k}M_{a}$, and $\operatorname{\mathbf{rk}}_{a,b}M:=\operatorname{\mathrm{rk}}M_{a\to b}$.
The split Grothendieck group $\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})$ of $\operatorname{\mathrm{mod}}\mathcal{P}$ is the free abelian group generated by isoclasses $[M]$ of objects $M$ in $\operatorname{\mathrm{mod}}\mathcal{P}$ modulo the relation $[M\oplus N]=[M]+[N]$ for any objects $M,N$ in $\operatorname{\mathrm{mod}}\mathcal{P}$. Fixing a set $\operatorname{\mathrm{Ind}}\mathcal{P}$ of representatives of isoclasses of the indecomposable $\mathcal{P}$-modules, the set $\{[U]\in\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\mid U\in\operatorname{\mathrm{Ind}}\mathcal{P}\}$ forms a basis for $\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})$. In other words, decomposing $M$ into indecomposable direct summands $M\simeq\bigoplus_{U\in\operatorname{\mathrm{Ind}}\mathcal{P}}U^{\alpha_{U}}$, the vector $(\alpha_{U})_{U\in\operatorname{\mathrm{Ind}}\mathcal{P}}$ uniquely determines $M$ up to isomorphism, by the Krull-Schmidt property.
@@
    For $a\in\mathcal{P}$, let $[a,\infty)=\{b\in\mathcal{P}\mid a\leq b\}$. The interval module $P_{a}:=\mathbb{I}_{[a,\infty)}$ is an indecomposable projective, and all indecomposable projective $\mathcal{P}$-modules have this form.
@@
    For $a\in\mathcal{P}$, let $(-\infty,a]=\{b\in\mathcal{P}\mid a\geq b\}$. The interval module $I_{a}:=\mathbb{I}_{(-\infty,a]}$ is an indecomposable injective. Note also that all indecomposable injective $\mathcal{P}$-modules have this form.
@@
    For $a\in\mathcal{P}$, the interval module $S_{a}:=\mathbb{I}_{\{a\}}$ (which is also a rectangle module since $\{a\}=[a,a]$) is the simple top of $P_{a}$ and the simple socle of $I_{a}$. Note that all simple $\mathcal{P}$-modules have this form.
@@
    For $a<b$, let $[a,b[\ =\{c\in\mathcal{P}\mid a\leq c\not\geq b\}$. The interval module $H_{a,b}:=\mathbb{I}_{[a,b[}$ is called a hook module, see [BOO24]. Note that $H_{a,b}$ is the cokernel of the natural injective map $P_{b}\rightarrow P_{a}$.
@@
    For $a<b$, let $]a,b]\ =\{c\in\mathcal{P}\mid b\geq c\not\leq a\}$. The interval module $C_{a,b}:=\mathbb{I}_{]a,b]}$ is called a cohook module. Note that $C_{a,b}$ is the kernel of the natural surjective map $I_{a}\rightarrow I_{b}$.

### v3

@@ Abstract.
One of the main objectives of topological data analysis is the study of discrete invariants for persistence modules, in particular when dealing with multiparameter persistence modules where complete invariants are unavailable. In many cases, the invariants studied for these non-totally ordered posets $\mathcal{P}$ can be obtained from restricting a given module $M\in\operatorname{\mathrm{mod}}\mathcal{P}$ to a subposet $\mathcal{X}$ of $\mathcal{P}$ that is totally ordered (or more generally, of finite representation type), and then computing the barcode (or the general direct sum decomposition) over $\mathcal{X}$.
In this paper, we systematically study invariants arising from order-preserving embeddings of representation-finite posets $\mathcal{X}$ into $\mathcal{P}$. Given such an embedding, we associate to a $\mathcal{P}$-module the multiplicities of indecomposable $\mathcal{X}$-modules appearing in its restriction, and we analyze the resulting invariants as the embedding varies. This framework recovers classical invariants such as the rank invariant, and allows for a unified treatment of several constructions appearing in the literature.
The restriction functor from $\operatorname{\mathrm{mod}}\mathcal{P}$ to $\operatorname{\mathrm{mod}}\mathcal{X}$ is known to be exact and thus admits both left and right adjoint functors, known as induction and co-induction functors. This allows us to obtain new homological insights, and in particular, we describe bases of the images of these invariants, thereby generalizing the notion of signed barcodes.
We also address the redundancy arising from considering all embeddings of a fixed poset by introducing families of embeddings of increasing size and restricting attention to selected indecomposable modules. This leads to more efficient descriptions of the invariants and yields new bases, including interval-module bases that generalize previously known constructions.
@@ 1.1. Setting
It turns out that, in many cases, these invariants studied for non-totally ordered posets $\mathcal{P}$ can be obtained from restricting a given module to a subposet $\mathcal{X}$ of $\mathcal{P}$ that is totally ordered (or more generally, of finite representation type), and then computing the barcode (or the general direct sum decomposition) over $\mathcal{X}$. This approach has been described in e.g. [AENY23b, DKM24, HNOX, La18, LW] introducing certain subposets $\mathcal{X}$ having the form of linear subposets, “zigzags”, or “tours” and “courses”, all referring to a poset whose indecomposable representations can be described via intervals. The full classification of those finite posets whose indecomposable representations can all be described via intervals can be found in [Cha69] and [AET25, Section 5].
@@
To fix notation, consider an order-embedding $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$; that is, a poset morphism that is an isomorphism onto its image. The restriction of a $\mathcal{P}$-module $M$ to $f(\mathcal{X})$ can be described completely by counting the multiplicities of indecomposable direct summands. We assume $\mathcal{X}$ to be representation-finite, and thus obtain a discrete invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}^{f}$ of rank $|\operatorname{\mathrm{Ind}}\mathcal{X}|$, the number of indecomposable $\mathcal{X}$-modules. Repeating this process for all possible embeddings $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$ and collecting the result into a vector indexed over all embeddings, we obtain the invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ describing all multiplicities of indecomposable $\mathcal{X}$-modules obtained from all possible ways to restrict the $\mathcal{P}$-module $M$ to a subposet of form $\mathcal{X}$. One also defines an equivalent invariant $\mathbf{dimh}_{\mathcal{X},\mathcal{P}}$ by calculating the Hom-dimensions from the set of indecomposable $\mathcal{X}$-modules to the restrictions of the persistence module along all order embeddings. See Definition 3.3 and Theorem 3.6 for details.
Consider for example the poset $\mathcal{X}_{1}$ with one element. Then $\mathcal{X}_{1}$ has only one indecomposable representation, and its multiplicity in a general module is given by the dimension. Thus listing all possible embeddings of $\mathcal{X}_{1}$ into $\mathcal{P}$ and counting multiplicities amounts to describing the dimension vector, so $\mathbf{mult}_{\mathcal{X}_{1},\mathcal{P}}=\operatorname{\mathbf{dim}}_{\mathcal{P}}$. Consider now the poset $\mathcal{X}_{2}=\{1\to 2\}$. This poset is representation-finite, and the multiplicity of the indecomposable representation given by the interval $[1,2]$ can be computed by calculating the rank of the map $M_{1}\to M_{2}$. This allows one to show that the multiplicity invariant $\mathbf{mult}_{\mathcal{X}_{2},\mathcal{P}}$ is equivalent to the well-studied rank invariant $\operatorname{\mathbf{rk}}_{\mathcal{P}}$, see Corollary 3.8. More generally, as discussed in Example 3.5, the “generalized rank invariant” of [KM21] and the “compressed multiplicities” of [AENY23b] both record the multiplicity of the sincere interval modules $\mathbb{I}_{\mathcal{X}}$ over certain subsets $\mathcal{X}\subseteq\mathcal{P}$ (see also [BDL25] for a more general discussion of the relation between these multiplicities and ranks).
@@ 1.2. Results
For several subposets $\mathcal{X}$, we address the question of how the multiplicities in $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ can be computed in terms of linear algebra. See Examples 3.7 and 3.15 and the calculations in Section 3. In each case, the multiplicities of indecomposable summands of $M$ can be obtained computing ranks or intersections given by the structure maps of the module $M$. More generally, we show in Theorem 3.6 that the invariant $\mathbf{mult}_{\mathcal{X},\mathcal{P}}$ can be obtained by calculating the Hom-dimensions from the set of indecomposable representations to the given persistence module, thus reducing the problem to solving a system of linear equations.
Besides computability, the second and equally important quality of an invariant for persistence modules is the stability. For the rank invariant, this question has been addressed in [La18], and further explored in [BOO24, BOOS24] where the authors single out a class of special persistence modules, the “hook modules”, which are defined as indicator modules of certain shapes called hooks: For $a<b$ in $\mathcal{P}$, let $a,b[\ =\{c\in\mathcal{P}\mid a\leq c\not\geq b\}$. The interval module $\mathbb{I}_{[a,b[}$ is called a hook module. Note that $\mathbb{I}_{[a,b[}$ is the cokernel of the natural map $\mathbb{I}_{[b,\infty)}\rightarrow\mathbb{I}_{[a,\infty)}$ between indecomposable projective $\mathcal{P}$-modules. It is shown in [BOO24] that the hook modules play the role of projective objects in a relative homology theory given by some exact structure $\mathcal{E}$, thus every persistence module $M$ admits a resolution by hook modules. One can then show that these hooks form a basis for the (image of the) rank invariant (see Definition 4.11), which in turn allows one to define a “signed barcode”, leading to a stability result for the considered invariant [BOOS24].
The viewpoint of relative homological algebra has further been extended in many recent works (see e.g. [AENY23a, BBH24, BBH, BOO24, BOOS24, CGRST24, OS24]), but we are only now able to conceptually explain the appearance and importance of hook modules and their generalizations: Studying embeddings $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$, the corresponding restriction functor $f^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{P}\to\operatorname{\mathrm{mod}}\mathcal{X}$ admits both a left adjoint $f_{!}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$ (called coinduction) and a right adjoint $f_{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{\mathrm{mod}}\mathcal{X}\to\operatorname{\mathrm{mod}}\mathcal{P}$ (called induction). One can then use the adjunction formula to show that the set
$\mathcal{H}=\{f_{!}U\mid U\mbox{ indecomposable }\mathcal{X}\mbox{-module and }f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}\mbox{ an embedding}\}$
@@
splits for all embeddings $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$. See Proposition 4.8. Coming back to the rank invariant, we study embeddings of the poset $\mathcal{X}_{2}$, and since every indecomposable $\mathcal{X}_{2}$-module is the cokernel of a map between indecomposable projectives, we see easily that the set $\mathcal{H}$ above is formed exactly by the hook modules in this case.
@@
Already in [BOO24], they specified a second basis for the rank invariant, given by rectangle modules $\mathbb{I}_{[a,b]}$ with support $[a,b]$, which are somewhat easier to deal with than the hook modules (note that one needs to consider right-open rectangles when the poset $\mathcal{P}$ is infinite). We aim to generalise this result here in terms of general embeddings: Given a poset embedding $f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\rightarrow\mathcal{P}$ and $U\in\operatorname{\mathrm{mod}}\mathcal{X}$, the adjunction formulas yield a natural map $\theta_{f}U\mathrel{\mathop{\ordinarycolon}}f_{!}U\rightarrow f_{*}U$. We denote $\Theta_{f}U=\mathrm{Im}(\theta_{f}U)$. (This is the so-called “intermediate extension functor”, see [AET25, Section 4.2] for further details and historical context.) One easily verifies that in case $\mathcal{X}=\mathcal{X}_{2}$, the set
[…]
@@
$\mathcal{H}=\{f_{!}U\mid U\mbox{ indecomposable }\mathcal{X}\mbox{-module and }f\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}\mbox{ an embedding}\}$
produces the same $\mathcal{P}$-modules several times from different embeddings $f$ and different indecomposables $U$, for instance the simple modules are double counted whenever the images of two different embeddings $f_{1},f_{2}\mathrel{\mathop{\ordinarycolon}}\mathcal{X}\to\mathcal{P}$ intersect. It is therefore more efficient to first collect all simple $\mathcal{P}$-modules (once) via embeddings of the one-point poset $\mathcal{X}_{1}$, then consider all two-dimensional $\mathcal{P}$-modules obtained from the interval $[1,2]$ by embedding the poset $\mathcal{X}_{2}=\{1\to 2\}$ into $\mathcal{P}$ and so on. In other words, instead of studying the set of all indecomposables obtained from embeddings of one poset $\mathcal{X}$, we study now embeddings of several posets of increasing sizes, but limit attention to only some indecomposables (that have not been obtained from embedding of smaller posets previously). The setting we consider here is where we fix a family of posets $\mathcal{F}=\{\mathcal{X}_{i}\}$ together with certain embeddings into $\mathcal{P}$, and then we define $\mathbf{mult}_{\mathcal{F}}$ to count the multiplicity of the characteristic indecomposable module $\mathbb{I}_{\mathcal{X}_{i}}$ in the restriction of $M$ to $\mathcal{X}_{i}$ along the various embeddings. So only one indecomposable is considered for each poset embedding. As discussed in Example 5.8, both the generalized rank invariant of [KM21] and the various compressed multiplicities of [AENY23b] can be realized in the form $\mathbf{mult}_{\mathcal{F}}$.
@@
With $\mathcal{F}\mathrel{\mathop{\ordinarycolon}}=\{\mathcal{X}_{1},\mathcal{X}_{2}\}$ for $\mathcal{X}_{1}$ and $\mathcal{X}_{2}$ the one- and two-point poset as before, we can recover the result that the set $\mathcal{R}$ (the rectangles in this case) forms a basis for $\mathbf{mult}_{\mathcal{F}}$.
Theorem 1.1.
@@
$\mathcal{R}=\{\Theta_{f}(U)\mid(f,U)\in\mathrm{Emb}_{\mathcal{X}}^{\mathcal{P}}\times\operatorname{\mathrm{Ind}}\mathcal{X}\}$
@@ 2. $\mathbb{Z}$-linear invariants
An object $M\in\operatorname{\mathrm{mod}}\mathcal{P}$ is also called a $\mathcal{P}$-module. We denote by $M_{a}$ the vector space $M(a)$, and by $M_{a\to b}$ the element in $\operatorname{\mathrm{Hom}}_{k}(M_{a},M_{b})$ given by the functor $M$ applied to the pair $a\leq b\in\mathcal{P}$ . We further denote $\operatorname{\mathbf{dim}}_{a}M\mathrel{\mathop{\ordinarycolon}}=\dim_{k}M_{a}$, and $\operatorname{\mathbf{rk}}_{a,b}M\mathrel{\mathop{\ordinarycolon}}=\operatorname{\mathrm{rk}}M_{a\to b}$.
The split Grothendieck group $\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})$ of $\operatorname{\mathrm{mod}}\mathcal{P}$ is the free abelian group generated by isoclasses $[M]$ of objects $M$ in $\operatorname{\mathrm{mod}}\mathcal{P}$ modulo the relation $M\oplus N]=[M]+[N]$ for any objects $M,N$ in $\operatorname{\mathrm{mod}}\mathcal{P}$. Fixing a set $\operatorname{\mathrm{Ind}}\mathcal{P}$ of representatives of isoclasses of the indecomposable $\mathcal{P}$-modules, the set $\{[U]\in\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})\mid U\in\operatorname{\mathrm{Ind}}\mathcal{P}\}$ forms a basis for $\mathrm{K}_{0}^{\mathrm{sp}}(\mathcal{P})$. In other words, decomposing $M$ into indecomposable direct summands $M\simeq\bigoplus_{U\in\operatorname{\mathrm{Ind}}\mathcal{P}}U^{\alpha_{U}}$, the vector $(\alpha_{U})_{U\in\operatorname{\mathrm{Ind}}\mathcal{P}}$ uniquely determines $M$ up to isomorphism, by the Krull-Schmidt property.
@@
    For $a\in\mathcal{P}$, let $a,\infty)=\{b\in\mathcal{P}\mid a\leq b\}$. The interval module $P_{a}\mathrel{\mathop{\ordinarycolon}}=\mathbb{I}_{[a,\infty)}$ is an indecomposable projective, and all indecomposable projective $\mathcal{P}$-modules have this form.
@@
    For $a\in\mathcal{P}$, let $(-\infty,a]=\{b\in\mathcal{P}\mid a\geq b\}$. The interval module $I_{a}\mathrel{\mathop{\ordinarycolon}}=\mathbb{I}_{(-\infty,a]}$ is an indecomposable injective. Note also that all indecomposable injective $\mathcal{P}$-modules have this form.
@@
    For $a\in\mathcal{P}$, the interval module $S_{a}\mathrel{\mathop{\ordinarycolon}}=\mathbb{I}_{\{a\}}$ (which is also a rectangle module since $\{a\}=[a,a]$) is the simple top of $P_{a}$ and the simple socle of $I_{a}$. Note that all simple $\mathcal{P}$-modules have this form.
@@
    For $a<b$, let $a,b[\ =\{c\in\mathcal{P}\mid a\leq c\not\geq b\}$. The interval module $H_{a,b}\mathrel{\mathop{\ordinarycolon}}=\mathbb{I}_{[a,b[}$ is called a hook module, see [BOO24]. Note that $H_{a,b}$ is the cokernel of the natural injective map $P_{b}\rightarrow P_{a}$.
@@
    For $a<b$, let $a,b]\ =\{c\in\mathcal{P}\mid b\geq c\not\leq a\}$. The interval module $C_{a,b}\mathrel{\mathop{\ordinarycolon}}=\mathbb{I}_{]a,b]}$ is called a cohook module. Note that $C_{a,b}$ is the kernel of the natural surjective map $I_{a}\rightarrow I_{b}$.